OpenStudy (anonymous):

What is the vertex of this quadratic function? f(x) = 2x2 + 4x − 5

OpenStudy (danjs):

which math class are you taking

OpenStudy (anonymous):

math2

OpenStudy (danjs):

algebra

OpenStudy (anonymous):

oh yeah

OpenStudy (danjs):

so you need to turn y = ax^2 + bx + c into y = a*(x-h)^2 + k the vertex form

OpenStudy (danjs):

the vertex in that is the point (h,k)

OpenStudy (anonymous):

i dont get it

OpenStudy (jenma):

U know how to find the axis of symmetry?

OpenStudy (danjs):

you can play with that , or simply remember, the x coordinate of the vertex of y = ax^2 + bx + c is x = -b / (2*a)

OpenStudy (danjs):

as jenma said , or also the axis of symm

OpenStudy (danjs):

then use that x value , and calculate what y is,, the point (x,y) is the vertex

OpenStudy (danjs):

here is an example i just found... http://hotmath.com/hotmath_help/topics/vertex-of-a-parabola.html

OpenStudy (danjs):

gl

OpenStudy (anonymous):

ty

OpenStudy (jenma):

Yes @DanJS . Find the axis of symmetry first and then when u find the answer u would do the substitute of the equation and u will find the vertex...